OpenStudy (anonymous):

4) Given the set of scores 5.4, 5.6, 5.7, 5.9, 6.0; find the standard deviation.

OpenStudy (anonymous):

average = \[\frac{ 5.4 + 5.6 + 5.7 + 5.9 + 6 }{ 5 } = 5.72\] variation = \[\frac{ (5.4 - 5.72)^2 + (5.6 - 5.72)^2 + (5.7 - 5.72)^2 + (5.9 - 5.72)^2 + (6 - 5.72)^2}{ 5 } = 0.0456\] standard deviation = \[\sqrt{variation} = \sqrt{0.0456} = 0.213542\]

OpenStudy (anonymous):

ok so what is the viriation

OpenStudy (anonymous):

ok I have another question 2) A local car dealership advertises cars that get the following miles per gallon:15, 22, 36, 30, 38, 21 Find the variance. (Round your answer to the nearest hundreth.) do you do it the same way

OpenStudy (anonymous):

Yes, variation is just how the numbers vary from the average. You just calculate the average of 15,22,36,30,38,21 and then \[\frac{(15-avg)^2+(22-avg)^2 ....}{5}\]Got it ? :)

OpenStudy (anonymous):

It should be divided by 6, sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

5) Eight golfers posted the following scores during the tournament: 78, 75, 82, 85, 74, 80, 80, 82 Find the standard deviation. (Round your answer to the nearest hundreth.)

OpenStudy (anonymous):

so far I know the average is 480.71

OpenStudy (anonymous):

and I did 78-480.71+75-480.71........ and I got -4440.81

OpenStudy (anonymous):

you have to square the numbers you substract so it should be \[(78-480.71)^2+(75-480.71)^2\] and so on

OpenStudy (anonymous):

and i got the answer - 4440.81

OpenStudy (anonymous):

what do i do now

OpenStudy (anonymous):

so you do the variation and the standard deviation the same way

OpenStudy (anonymous):

k first of all the average is not 480.71 but 79.5 (You sum up all the scores and divide by the number of scores you have. After that you calculate the variation as follows \[\frac{(78-79.5)^2+(75-79.5)^2+(82-79.5)^2+(85-79.5)^2+(74-79.5)^2+(80-79.5)^2+(80-79.5)^2+(82-79.5)^2}{8}\] after that you take the square root of variation which is the standard deviation \[\sqrt{12}=3.46\] done

OpenStudy (anonymous):

so you do the variation and the standard deviation the same way

OpenStudy (anonymous):

i did 74+75+78+80+80+82+82/7 and that's how i got the average

OpenStudy (anonymous):

i got 23.9622098313156 as the answer